# Cyber-physical WebAssembly:

### v1

@@ V Implementation Details
To demonstrate the feasibility of the proposed cyber-physical WebAssembly architecture, we developed several proof-of-concept implementations, each addressing different aspects of the proposed solution. Our primary objectives were to validate the proposed approach and identify potential opportunities or shortcomings in the underlying architecture and standard methodologies. In particular, this research presents three distinct proof-of-concept implementations.
@@
  • I2C in Wasmtime: Demonstrates driving an I2C seven-segment display from WebAssembly running on a Linux host and on a Raspberry Pi Pico microcontroller, as well as reading temperature and humidity data from an HTS221 sensor on a Raspberry Pi 3B.
@@
  • I2C in WAMR: Provides a proof of concept that can read data from an HTS221 sensor. Note that this solution does not yet use the component model or capability-based security due to the current limitations of WAMR, including the lack of support for WASI Preview 2. Instead, an ACL-based Security System was implemented.
@@
  • USB in Wasmtime: Showcases the communication of WebAssembly applications with USB devices, including a USB flash drive using a WebAssembly-based driver and an Xbox controller using a separate driver.
The remainder of this section describes each proof-of-concept implementation.
V-A I2C in Wasmtime
@@
The host implementation handles the actual I2C connection with the device, whereas the guest implementations are responsible for device-specific logic, such as reading the temperature or writing the time to the display. The host and guest components interact with each other through two WIT worlds, one for the seven-segment display and the other for the HTS221 sensor. These WIT worlds import the necessary interfaces from the WASI-I2C proposal that are implemented by the host, allowing the guest to communicate with the I2C device. Additionally, these WIT worlds export device-specific functions, such as get-temperature, that are implemented and made available by the guest. Although these worlds were written for our specific proof of concepts, they aim to be as generic as possible to allow future support for other displays and sensors. For example, the display’s write function can accept any string, even though our display only has four seven-segment characters.
@@
The guest implementations do not make use of Rust’s std library because they only need specific types that can be imported individually, which also helps minimize the size of the compiled Wasm binaries. The guest implementation for the seven-segment display directly writes the necessary bytes to control it, while the guest component for the HTS221 sensor leverages the embedded-hal compatible hts221 [37] Rust crate for simplicity. To adapt the embedded-hal API to the slightly different generated bindings, the wasi-embedded-hal [38] crate was developed.
V-B I2C in WAMR
@@
WAMR currently does not support WASI Preview 2. Consequently, the component model and WIT cannot be used in WAMR, which means that a standardized I2C connection cannot be passed to the guest because WIT resources are not available. As a workaround, the I2C connection is kept global on the host side. For the host to be able to bind with the guest, it has to register the function that the guest imports and find the functions that the guest exports. Parameters are passed to the guest’s exported functions using a vector of WasmValue. Internally, a WasmValue is represented as an array of 32-bit unsigned integers.
The ACL security system for the I2C implementation employs a struct i2c_context to store access permissions. A reference is registered in the user_data field of the execution environment, which contains the device address, indicating that the specific module is permitted to communicate with the device. As it is necessary to reestablish the I2C connection on each I2C API call, a check for permissions can be performed. Given the frequency with which I2C functions are called in embedded applications, it is necessary to ensure that the lookup time is constant. To this end, the allowed addresses are stored in two 64-bit fields spanning the full range of 128 valid I2C addresses, thereby ensuring that the lookup time remains consistent.
V-C USB in Wasmtime
@@
The proof-of-concept implementation [39] of USB in Wasmtime includes a host implementation along with several guest implementations, such as interfacing with a USB flash drive or an Xbox controller. The main concept of the USB implementation in Wasmtime mirrors that of the I2C implementation in Wasmtime: the host component is platform dependent and interacts with the underlying system to access the USB resource, while the guest components handle device-specific logic, such as reading the joystick positions of an Xbox controller.
@@
To facilitate the development of the host component and enhance its platform independence, rusb [40], a Rust wrapper around libusb [32], is used. The USB flash drive guest component implements a range of commands for USB Attached SCSI devices, including basic file system operations like listing directory contents and reading files. To streamline the proof-of-concept implementation, it utilizes the fatfs [41] Rust crate for file system support. The file system support is fully integrated within the guest component itself, without relying on the host operating system’s file system capabilities. The Xbox guest component continuously monitors the current state of the controller through its USB interrupt pipe and writes button presses and joystick positions to standard output.
VI Evaluation Results
@@
VI-A I2C Performance Evaluation
The I2C proof of concept for the seven-segment display is evaluated on a Raspberry Pi 4B, whereas the HTS221 sensor proof of concept is evaluated on a Raspberry Pi 3B. The display evaluation will compare the native, Wasmtime, and WAMR implementations, while the HTS221 sensor evaluation will compare the native and Wasmtime implementations. All benchmark results are available on GitHub [36].
To assess performance overhead, we conduct 100 iterations of writing a four-digit number to the display and also reading the temperature from the HTS221 sensor. To highlight the worst-case overhead, cold start execution times are measured, where each measurement includes the time required to instantiate the Wasm binary within the runtime. The mean cold start execution times are then compared across the native, Wasmtime, and WAMR implementations. Fig. 2 illustrates the mean cold start execution times for the two proof of concepts. Wasmtime shows a notable overhead relative to the native implementation, with the mean cold start execution times being 7.16 ms (+238.7%) longer for the display and 1.38 ms (+18.9%) longer for the HTS221 proof of concept. In contrast, the WAMR implementation incurs only a minimal overhead of 50 $\mu$s (+1.7%). Most of the additional time consumed by Wasmtime can be attributed to functions associated with Cranelift [42], a compiler backend utilized by Wasmtime for both just-in-time and ahead-of-time (AOT) compilation.
@@ VI-A I2C Performance Evaluation
VI-B USB Performance Evaluation
The USB proof of concept is evaluated on three systems: one with an AMD Ryzen 7 5800H and USB3.0 running x86 Linux 6.8.9, a Raspberry Pi 3B+ with USB2.0 running AArch64 Linux 6.8.0, and a third with an AMD Ryzen 7 5800X3D and USB3.0 running x86 Windows 10 22H2. All benchmark results are available on GitHub [43].
To assess the runtime size and memory usage overhead introduced by the USB host component, we compare the Wasmtime runtime both with and without the USB host component. Our measurements show that the runtime size overhead is minimal, adding only 160 KiB (+1.1%) for the x86 Linux runtime, 220 KiB (+1.7%) for the AArch64 Linux runtime, and 275 KiB (+1.9%) for the Windows runtime. The memory usage overhead is also minimal for the Linux runtimes, comprising only 459 KiB (+1.7%) for the x86 Linux runtime and 295 KiB (+1.5%) for the AArch64 Linux runtime. The Windows version of the runtime incurs a more severe memory overhead of 1427 KiB (+11.0%), which can be attributed to an entirely different backend being used by libusb on Windows.
@@ VI-B USB Performance Evaluation
The latency overhead is evaluated using an Arduino Nano 33 IoT device with custom firmware that receives a message from the host device and echoes it back to the host without modifications. This allows the host to measure the round-trip time (RTT) of this message. Fig. 4 shows the observed RTTs for a bulk transfer of a 32-byte message. Due to the low priority of bulk transfers among USB transfer modes, numerous outliers are present. Some outliers have been excluded to facilitate the comparison of the median RTTs. The overhead introduced by WebAssembly is relatively modest. On x86 Linux, it results in a median RTT that is only 7 $\mu$s (+3.5%) longer than the native implementation. On x86 Windows, an even lower overhead of 2 $\mu$s (+0.8%) is observed. However, on AArch64, the overhead is more pronounced, with an increase of 45 $\mu$s (+8.0%).
@@
VII Conclusion
With the growing popularity of WebAssembly on embedded and IoT devices, the need for standardized hardware interface support in WebAssembly arose. This paper introduces two WebAssembly System Interface proposals, WASI-I2C and WASI-USB, enabling the use of I2C and USB hardware interfaces within WebAssembly. The cyber-physical WebAssembly proof-of-concept architecture is described in depth, including a solution for authorized hardware interface access. Hardware interface access is managed through the capability-based security mechanism of WASI, and is extended with explicit ACL-based security for more granular access control. To ensure our proposals remain compatible with future toolchains and platforms, we have initiated the standardization process in collaboration with industry partners. Development details for the standardized proposals are provided, highlighting key decisions made throughout the process. Proof-of-concept implementations showcase the feasibility of the proposed architecture and are evaluated to confirm that they meet functional requirements and to identify any performance impacts. Our evaluation results show that the overhead introduced by WASI-USB is minimal, resulting in a maximum throughput reduction of 1.4 MiB/s (-4.1%) and an increase in round-trip time of up to 45 $\mu$s (+8.0%). Our test results for WASI-I2C include the initialization of the Wasm binary within the runtime in each measurement. This reveals significantly worse overhead in Wasmtime, with the mean cold start execution time increasing by up to 7.16 ms (+238.7%) at worst and 1.38 ms (+18.9%) at best. This emphasizes that the overhead of runtime and binary initialization is significant, indicating that low-latency applications require a warm-up period.
@@ Acknowledgment
We extend our heartfelt gratitude to everyone who contributed to these proposals: Valentin Olpp, Dan Gohman, Emiel Van Severen and the Bytecode Alliance & WASI subgroup of the W3C WebAssembly Community Group.
This work received funding from the Smart Networks and Services Joint Undertaking (SNS JU) under the Horizon Europe research and innovation program of the European Union under Grant Agreement number 101139067 (https://elasticproject.eu/). Expressed views and opinions are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ IV Implementation Details
To validate our proposed cyber-physical WebAssembly architecture, we developed three proof-of-concept implementations, each addressing different aspects of the solution. Our primary objectives were to confirm the feasibility of the approach and identify potential opportunities or shortcomings in the underlying architecture and standardization methodologies.
@@
  • I2C in Wasmtime: Demonstrates driving an I2C seven-segment display from Wasm on a Linux host and a Raspberry Pi Pico, and reading temperature and humidity data from an HTS221 sensor on a Raspberry Pi 3B.
@@
  • I2C in WAMR: Demonstrates reading data from an HTS221 sensor. Due to WAMR’s current limitations, including the lack of WASI Preview 2 support, this solution uses an ACL-based security system instead of the component model or capability-based security.
@@
  • USB in Wasmtime: Demonstrates Wasm applications communicating with USB devices, including a flash drive and an Xbox controller using Wasm-based drivers.
IV-A I2C in Wasmtime
@@
The host implementation handles the I2C connection with the device, whereas the guests are responsible for device-specific logic, such as reading temperature or displaying time. The host and guest components interact through two WIT worlds, one for the display and the other for the HTS221 sensor. These WIT worlds import the necessary interfaces from the WASI-I2C [27] proposal that are implemented by the host, allowing the guest to communicate with the I2C device. Additionally, the WIT worlds export device-specific functions implemented by the guest. Although these worlds were written for our specific proof of concepts, they aim to be as generic as possible to allow future support for other displays and sensors.
@@
The guest implementations do not use Rust’s std library, only importing necessary types individually, which helps minimize the size of the compiled Wasm binaries. The guest component for the seven-segment display directly writes bytes to control it, while the HTS221 sensor component leverages the embedded-hal [28] compatible hts221 [29] Rust crate for simplicity. To adapt the embedded-hal API to the slightly different generated bindings, the wasi-embedded-hal [30] crate was developed.
IV-B I2C in WAMR
@@
WAMR currently does not support WASI Preview 2. Consequently, the component model and WIT are not available, which means that a standardized I2C connection cannot be passed to the guest. As a workaround, the I2C connection remains global on the host side. For the host to bind with the guest, it has to register the function that the guest imports and find the functions that the guest exports.
IV-C USB in Wasmtime
@@
The proof-of-concept implementation [31, 32] of USB in Wasmtime includes a host implementation along with several guest implementations. The main concept of the USB implementation in Wasmtime mirrors that of the I2C implementation in Wasmtime: the host component is platform dependent and interacts with the underlying system to access the USB resource, while the guest components handle device-specific logic, such as reading the inputs of an Xbox controller.
@@
To facilitate the development of the host component and enhance its platform independence, rusb [33], a Rust wrapper around libusb [34], is used. The USB flash drive guest component supports various USB Attached SCSI device commands, including basic file system operations like listing directory contents and reading files. The file system support is fully integrated within the guest component itself using the fatfs [35] Rust crate, without relying on the host operating system’s file system capabilities. The Xbox guest component continuously monitors the current state of the controller through its USB interrupt pipe and writes button presses and joystick positions to standard output.
V Evaluation Results
@@
V-A I2C Performance Evaluation
The I2C proof of concept for the seven-segment display is evaluated on a Raspberry Pi 4B, whereas the HTS221 sensor proof of concept is evaluated on a Raspberry Pi 3B. The display evaluation will compare the native, Wasmtime, and WAMR implementations, while the HTS221 sensor evaluation will compare the native and Wasmtime implementations. All benchmark results are available [26] on GitHub.
To assess performance overhead, we conduct 100 iterations of writing a four-digit number to the display and also reading the temperature from the HTS221 sensor. To highlight the worst-case overhead, cold start execution times are measured, where each measurement includes the time required to instantiate the Wasm binary within the runtime. The mean cold start execution times are then compared across the native, Wasmtime, and WAMR implementations. Fig. 2 illustrates the mean cold start execution times for the two proof of concepts. Wasmtime shows a notable overhead relative to the native implementation, with the mean cold start execution times being 7.16 ms (+238.7%) longer for the display and 1.38 ms (+18.9%) longer for the HTS221 proof of concept. In contrast, the WAMR implementation incurs only a minimal overhead of 50 $\mu$s (+1.7%). Most of the additional time consumed by Wasmtime can be attributed to functions associated with Cranelift [36], a compiler backend utilized by Wasmtime for both just-in-time and ahead-of-time compilation.
@@ V-A I2C Performance Evaluation
V-B USB Performance Evaluation
The USB proof of concept is evaluated on three systems: one with an AMD Ryzen 7 5800H and USB3.0 running x86 Linux 6.8.9, a Raspberry Pi 3B+ with USB2.0 running AArch64 Linux 6.8.0, and a third with an AMD Ryzen 7 5800X3D and USB3.0 running x86 Windows 10 22H2. All benchmark results are available [37] on GitHub.
To assess the runtime size and memory usage overhead introduced by the USB host component, we compare the Wasmtime runtime both with and without the USB host component. Our measurements show that the runtime size overhead is minimal, adding only 160 KiB (+1.1%) for the x86 Linux runtime, 220 KiB (+1.7%) for the AArch64 Linux runtime, and 275 KiB (+1.9%) for the Windows runtime. Memory usage overhead is also minimal for the Linux runtimes, comprising only 459 KiB (+1.7%) for the x86 Linux runtime and 295 KiB (+1.5%) for the AArch64 Linux runtime. The Windows runtime incurs a more severe memory overhead of 1427 KiB (+11.0%), which can be attributed to an entirely different backend being used by libusb on Windows.
@@ V-B USB Performance Evaluation
The latency overhead is evaluated using an Arduino Nano 33 IoT device with custom firmware that receives a message from the host device and echoes it back to the host without modifications. This allows the host to measure the round-trip time (RTT) of this message. Fig. 4 shows the observed RTTs for a bulk transfer of a 32-byte message. Due to the low priority of bulk transfers among USB transfer modes, numerous outliers are present. Some outliers have been excluded to facilitate the comparison of the median RTTs. The overhead introduced by Wasm is relatively modest. On x86 Linux, it results in a median RTT that is only 7 $\mu$s (+3.5%) longer than the native implementation. On x86 Windows, an even lower overhead of 2 $\mu$s (+0.8%) is observed. However, on AArch64, the overhead is more pronounced, with an increase of 45 $\mu$s (+8.0%).
@@
VI Conclusion
With the growing popularity of WebAssembly on embedded and IoT devices, the need for standardized hardware interface support in WebAssembly arose. This paper introduces two WebAssembly System Interface proposals, WASI-I2C and WASI-USB, enabling the use of I2C and USB within WebAssembly. Our cyber-physical WebAssembly proof-of-concept architecture is described in depth, including a solution for authorized hardware interface access. Hardware interface access is managed through the capability-based security mechanism of WASI, and is extended with explicit ACL-based security for more granular access control. Proof-of-concept implementations showcase the feasibility of the proposed architecture and are evaluated to confirm that they meet functional requirements and to identify any performance impacts. Our evaluation results show that the overhead introduced by WASI-USB is minimal, resulting in a maximum throughput reduction of 1.4 MiB/s (-4.1%) and an increase in round-trip time of up to 45 $\mu$s (+8.0%). Our test results for WASI-I2C include the initialization of the Wasm binary within the runtime in each measurement. This reveals significantly worse overhead in Wasmtime, with the mean cold start execution time increasing by up to 7.16 ms (+238.7%) at worst and 1.38 ms (+18.9%) at best. This emphasizes that the overhead of runtime and binary initialization is significant, indicating that low-latency applications require a warm-up period.
@@ Acknowledgment
We extend our heartfelt gratitude to everyone who contributed to these standardization proposals: Valentin Olpp, Dan Gohman, Emiel Van Severen and the Bytecode Alliance & WASI subgroup of the W3C WebAssembly Community Group. This work received funding from the Smart Networks and Services Joint Undertaking (SNS JU) under the Horizon Europe research and innovation program of the European Union under Grant Agreement number 101139067 (https://elasticproject.eu/). Expressed views and opinions are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.